# Sequential Quadratic Sum-of-squares Programming for Nonlinear Control Systems

### v1

@@ IV-B Feasibility Restoration
where $\rho_{k}>0$ is a penalty parameter, and $\xi^{k}$ is the iterate at which the restoration phase is initiated. Problem (18) is similar to (14), but it differs in that both the decision variables and the constraint violations are optimized simultaneously. The additional decision variables introduced are negligible compared to the overall problem size. Due to the explicit reduction of constraint violation in the cost function, previous approaches for nonlinear parameter optimization, such as [51], needed a smooth reformulation of the problem using slack variables. In contrast, our approach is already smooth.
@@
Problem (18) is again nonlinear and hence we apply the previously described filter line-search algorithm to solve this problem. Two modifications are made:
@@
  • the restoration phase is left once an iterate is acceptable to the filter of the original problem with an additional threshold on constraint violation, which takes the form $\theta(\xi_{R}^{k})\leq\eta\theta(\xi^{k})$ with $\eta\in[0,1]$ (cf. [51]).
@@ 3 Solve (3) to obtain $\xi_{R}^{*}$ and $\lambda_{R}^{*}$ and set $\alpha_{k}=1$
5      Execute filter line-search (Algorithm 2)
@@ IV-C Implementation Details
IV-C1 Hessian Approximation
In order to obtain an accurate but also positive definite estimate of the Hessian, several methods can be found in the literature. Our implementation provides several options of which the applicability depends, e.g., on the specific problem size. Options include the damped Broyden-Fletcher-Goldfarb-Shanno (BFGS) update [36, Procedure 18.2] and exact Hessian computation in combination with regularization techniques such as the Gershgorin bound [5, Section 2.5], minimum Frobenius norm approach [36, p. 50] or using the mirrored version of the exact Hessian [43]. The initialization of the Hessian is also crucial for convergence of the algorithm. The matrix $H_{0}$ in Alg. 3 is initialized with the identity matrix, unless the user provides a scaling factor to improve convergence (cf. [36, Ch. 6]) or the exact (regularized) Hessian is used.
IV-C2 Second-Order Correction
Merit functions or filters may converge slowly because large steps are often rejected, a phenomenon known as the Maratos effect. A potential cause of slow convergence is an inaccurate linear approximation of the constraint function. To mitigate this, a second-order correction (SOC) step is applied to adjust the search direction and further reduce constraint violations [36]. Quadratic approximations of the constraint functions would address this but lead to a challenging quadratically-constrained quadratic program. An alternative, outlined in [36, p. 544], is employed here. The approach involves replacing the quadratic term in the constraint function’s quadratic approximation with a constant quadratic term based on the current uncorrected search direction $\omega^{k}=(\xi_{+}-\xi^{k})$ (see [36, p. 544] for details). Specifically, the constant term
@@ IV-C2 Second-Order Correction
IV-C3 Scaled Termination and Parameter
To improve robustness against numerical errors and poor scaling, we make use of a scaled KKT condition given by [44]
@@ IV-C3 Scaled Termination and Parameter
For a derivation of the scaled KKT condition see [35]. In our implementation, we set the tolerances and parameters of the filter as follows: $\epsilon_{\text{opt}}=10^{-4}$, $\epsilon_{\text{feas}}=10^{-6}$, $s_{f}=2$, $s_{\theta}=0.9$, $\delta=1$ and $\rho=10^{-4}$, $\gamma_{f}=\gamma_{\theta}=10^{-5}$, $\eta=10^{-4}$, $\rho_{\mathrm{min}}=0.01$ and $\rho_{\mathrm{max}}=1$.
V Case Studies
In this section we provide several numerical examples known from nonlinear system analysis and control design. Only nonconvex SOS problems with potentially non-affine system dynamics are considered. The proposed filter line-search algorithm is benchmarked to the state-of-the art method, that is, coordinate-descent (CD) (cf.  [8, 9, 24, 46, 52, 53, 26, 11, 33, 30, 34, 3, 10, 48, 45, 32]). We compare the number of iterations, computation time and cost for given problem size. We denote the problem size by the number of system states $n$, constraints $n_{\text{con.}}$ (both linear and cone constraints) and decision variables $n_{\text{dec.}}$ of the underlying SDP. We report only the problem size of the sequential SOS approach, as it reflects the overall size. In contrast, CD involves multiple smaller subproblems, which may include redundant constraints or variables. A detailed breakdown is available in the supplementary material [39]. This section concludes with a comparison of the main characteristics of the two methods and provides practical aspects such as initial guess generation for the sequential approach.
V-A Preliminaries
[…]
V-B Region-of-Attraction Estimation
We consider the problem to find the region-of-attraction (ROA) for nonlinear closed-loop systems. We are interested to solve the following standard ROA estimation problem
@@ V-B Region-of-Attraction Estimation
where $V$ is the sought Lyapunov function, $s$ is a SOS multiplier coming from the generalized $\mathcal{S}$-procedure and $f(\cdot)$ is the closed loop dynamic, $g_{0}$ is a pre-defined safe set which we make use of to find a less conservative solution by minimizing the squared $l_{2}$-norm $\|\cdot\|_{\mathbb{R}[x]}$ on the space of polynomials. [^1] A small cost value indicates large portions of the safe set are covered by the ROA function. The stable level set is pre-defined as $\gamma=1$, but could be also a decision variable. Two examples are considered: The first example is taken from [9], where the nonlinear ROA analysis is used to assess the robustness properties of the F/A-18 flight control laws. The second example evaluates scalability and robustness in terms of the number of system states $n$, using an $N$-link (planar) robot arm.
V-B1 F/A-18
We take the seven state closed-loop dynamics from [9, Appendix B]. For both methods, we compute an initial Lyapunov candidate by linearizing the nonlinear closed-loop dynamics. For the sequential approach, we additionally set the initial SOS multiplier $s=x^{\top}x$. Due to infeasibility of CD at the first iteration, we solved the first subproblem via bisection. After the first iteration, we used the pre-defined level set. The results of the two approaches are summarized in Table II.
The sequential approach only needs 19 iterations until convergence, whereas coordinate descent did not converge within the pre-assigned maximum number of iterations. The total time (solver build + solve time) as stated in Table II is substantially less for sequential. The build time for the sequential SOS is $4.81\text{\,}\mathrm{s}$, whereas for coordinate descent, which includes several subproblems, is only $1.15\text{\,}\mathrm{s}$.
TABLE II: Comparison of sequential SOS (S) and coordinate descent (C) for the F-18 ROA problem.
@@ V-B1 F/A-18
V-B2 N-link Robot Arm
Using an $N$-link (planar) robot arm, the problem scales to $n=2N$ states. For each $N$, the dynamics and a linear quadratic control law (LQR) are computed and a quadratic polynomial approximation for the closed-loop system is calculated. The details are provided in the supplementary material [39]. We solve the problem twice using different initial guesses. In the first case, both algorithms are initialized with the Lyapunov function of the linear system, with the sequential algorithm additionally using a simple quadratic polynomial for the SOS multiplier. In the second, we make use of a clearly infeasible initial guess. A more detailed breakdown is given in Table III and  IV.
Good Initial Guess
The results are summarized in Table III. Both methods achieve optimality (using the assigned thresholds). Final costs are omitted because they are nearly identical, with the sequential method performing slightly better. Notably, the sequential method requires significantly less iterations than coordinate descent—especially for larger number of states—resulting in lower total computation time as shown in Fig. 2 and Table III. In some cases, CD was infeasible in the first iteration in the multiplier subproblem. In these cases we gave CD a grace iteration as indicated by and tried to solve the Lyapunov step. On average, sequential requires 60% less iterations and 54% less time to solve the problems.
Bad Initial Guess
The second case considers the case with a bad initial guess. We initialize all decision variables with negative quadratic polynomials. Such polynomials are clearly infeasible.
Remark 2.
It is unlikely that someone uses such an obviously infeasible initial guess and rather uses an initial guess similar to the first case. This is only made for demonstration purpose. It should be noted that an initial guess like in the first case does not guarantee convergence for coordinate descent in other cases, as demonstrated in the previous F/A-18 example.
Coordinate descent is not able to start the computation because it needs a feasible initial guess. Thus, no results can be reported in this case. In contrast, the proposed sequential algorithm with its feasibility restoration is able to solve the problem to full optimality or in a few cases at least to a feasible solution (indicated by $(\cdot)^{*}$ in Table IV). [^2] Feasible solution means the thresholds for constraint violation are met, but solution might be far from optimal. This clearly shows the enhanced robustness of the proposed approach. The call of the feasibility restoration phase increases the number of iterations significantly for smaller number of states. For medium to large number of states, the number of iterations is only moderately increased. However, the average time per iteration is much larger compared to cases where the feasibility restoration is not called, which is expected.
[…]
Figure 2: Relative performance improvement of sequential SOS (S) with respect to coordinate-descent (CD) in terms of total computation time $T_{(\cdot)}$ and iterations $I_{(\cdot)}$.
TABLE III: Comparison of sequential SOS (S) and coordinate descent (C) for the N-link robotic arm ROA example using a good initial guess.
n Size Iter. [-] Time [s]
$_{\text{con.}} n$$_{\text{dec.}}$ S C S C
@@ V-B2 N-link Robot Arm
TABLE IV: Comparison of sequential SOS (S) and coordinate descent (C) for the N-link pendulum ROA example using a bad initial guess.
@@
$_{\text{con.}} n$$_{\text{dec.}}$ S C S C
44 $85$ $248$ $45$ NaN $3.6$ NaN
66 $245$ $843$ $56$ NaN $7.8$ NaN
88 $558$ $2136$ $73$ NaN $30.9$ NaN
$10$ $1100$ $4535$ $11$ NaN $16.7$ NaN
$12$ $1963$ $8544$ $12$ NaN $52.3$ NaN
$14$ $3255$ 14 76314\,763 $11$ NaN $150.2$ NaN
$16$ $5100$ 23 88823\,888 $12$ NaN $454.2$ NaN
$18$ $7638$ 36 71136\,711 $16$ NaN $1354.01$ NaN
2020 11 02511\,025 54 12054\,120 $15$ NaN $3073.6$ NaN
$22$ 15 43315\,433 77 09977\,099 $14$ NaN $9873.7$ NaN
$24$ 21 05021\,050 106 728106\,728 $15$ NaN 22 245.922\,245.9 NaN
V-C Control Synthesis for Nonlinear Systems
The second case considers the synthesis of a stabilizing control law for nonlinear systems which consider state and control constraints. Besides of directly using such stabilizing control laws, such problems are relevant, for example, to compute terminal conditions needed to ensure asymptotic stability of nonlinear model predictive control. We consider the input constraint set as polyhedral set
@@ V-C Control Synthesis for Nonlinear Systems
with $H_{\mathcal{U}}\in\mathbb{R}^{p\times m}$, $p\in\mathbb{N}$, and $\mathbf{1}_{p}=(1,..,1)\in R^{p}$.
We are interested to solve the following SOS problem
@@
where $V(\cdot)$ is a the sought Lyapunov function, $\beta>0$ is the level set (either fixed or an additional decision variable), $\hat{\kappa}$ is the sought polynomial control law, and $s(\cdot)$ are the SOS multiplier resulting from Theorem 1. Equation (24b) and (24c) are standard Lyapunov conditions and (24d) and (24e) enforce state and control constraints, respectively. We make use of a similar cost function as before.
Again two examples are considered: The first example considers non-affine nonlinear dynamics of an aircraft’s longitudinal motion with four states and two controls, which cannot be solved with coordinate-descent in a direct manner. The second showcases the synthesis of a control law for control-affine satellite attitude dynamics.
V-C1 Nonlinear Aircraft Dynamics
In this example, we make use of the four state longitudinal motion of an aircraft provided in [8]. Compared to the previous examples, the control enter non-affine into the dynamics and hence, can not be solved by coordinate descent in a direct manner. For synthesis, we must re-write the system dynamic into an affine form. This would be possible by augmenting the system dynamics with a zero dynamic of the controls, resulting in $n_{e}=n+m$ state and hence in a higher computational load. In contrast, with the sequential approach we are able to directly solve this problem. In Table V, we provide only results for the sequential SOS approach. In about $13\text{\,}\mathrm{s}$ one obtains a safe-by-design controller for the nonlinear system.
V-C2 Satellite Attitude Control Synthesis
[…]
TABLE V: Comparison of sequential SOS (S) and coordinate descent (C) for the four-state constrained ROA problem.

### v2

@@ 4.3 Feasibility Restoration
where $\rho_{k}>0$ is a fixed penalty parameter, and $\xi^{k}$ is the iterate at which the restoration phase is initiated. Problem (19) is similar to (15), but it differs in that both the decision variables and the constraint violations are optimized simultaneously. The additional decision variables introduced are negligible compared to the overall problem size. Due to the explicit reduction of constraint violation in the cost function, previous approaches for nonlinear parameter optimization, such as wachterImplementationInteriorpointFilter2006a , needed a smooth reformulation of the problem using slack variables. In contrast, our approach is already smooth.
@@
Problem (19) is again nonlinear and hence we apply the previously described filter line search algorithm to solve this problem. Two modifications are made:
  1. 1.
@@
  2. 2.
    the restoration phase is left once an iterate is acceptable to the filter of the original problem with an additional threshold on constraint violation, which takes the form $\theta(\xi_{R}^{k})\leq\eta\theta(\xi^{k})$ with $\eta\in[0,1]$ (cf. wachterImplementationInteriorpointFilter2006a ).
@@ 3 Solve (4) to obtain $\xi_{R}^{*}$ and $\lambda_{R}^{*}$ and set $\alpha_{k}=1$
5      Execute filter line search (Algorithm 2)
@@ 4.4 Implementation Details
4.4.1 Hessian Approximation
In order to obtain an accurate but also positive definite estimate of the Hessian, several methods can be found in the literature. Our implementation provides several options of which the applicability depends, e.g., on the specific problem size. Options include the damped Broyden-Fletcher-Goldfarb-Shanno (BFGS) update (nocedalNumericalOptimization2006, , Procedure 18.2) and exact Hessian computation in combination with regularization techniques such as the Gershgorin bound (betts2010, , Section 2.5), minimum Frobenius norm approach (nocedalNumericalOptimization2006, , p. 50) or using the mirrored version of the exact Hessian quirynen2014 . The initialization of the Hessian is also crucial for convergence of the algorithm. The matrix $H_{0}$ in Alg. 3 is initialized with the identity matrix, unless the user provides a scaling factor to improve convergence (cf. (nocedalNumericalOptimization2006, , Chapter 6)) or the exact (regularized) Hessian is used.
4.4.2 Second-Order Correction
Merit functions or filters may converge slowly because large steps are often rejected, a phenomenon known as the Maratos effect. A potential cause of slow convergence is an inaccurate linear approximation of the constraint function. To mitigate this, a second-order correction (SOC) step is applied to adjust the search direction and further reduce constraint violations nocedalNumericalOptimization2006 . Quadratic approximations of the constraint functions would address this but lead to a challenging quadratically-constrained quadratic program. An alternative, outlined in (nocedalNumericalOptimization2006, , p. 544), is employed here. The approach involves replacing the quadratic term in the constraint function’s quadratic approximation with a constant quadratic term based on the current uncorrected search direction $\omega^{k}=(\xi_{+}-\xi^{k})$ (see (nocedalNumericalOptimization2006, , p. 544) for details). Specifically, the constant term
@@ 4.4.2 Second-Order Correction
4.4.3 Scaled Termination and Parameter
To improve robustness against numerical errors and poor scaling, we make use of a scaled KKT condition given by rauVski
@@ 4.4.3 Scaled Termination and Parameter
For a derivation of the scaled KKT condition see nikolayzik2012 . In our implementation, we set the tolerances and parameters of the filter as follows: $\epsilon_{\text{opt}}=10^{-4}$, $\epsilon_{\text{feas}}=10^{-6}$, $s_{f}=2$, $s_{\theta}=0.9$, $\delta=1$ and $\rho=10^{-4}$, $\gamma_{f}=\gamma_{\theta}=10^{-5}$, $\eta=10^{-4}$, $\rho_{\mathrm{min}}=0.01$ and $\rho_{\mathrm{max}}=1$.
5 Benchmarks
We provide several numerical benchmarks arising from nonlinear system analysis and control design. Only nonconvex SOS problems with potentially nonaffine system dynamics are considered. The proposed filter line search algorithm is benchmarked against the hitherto state-of-the-art method, coordinate-descent (CD, cf.  chakrabortyEtAl2011 ; chakrabortyEtAl2011ii ; iannelli2018 ; seilerBalas2010 ; yinEtAl2019 ; Yin2021 ; jarvisEtAl2003 ; Cunis2021aut ; majumdarEtAl2013 ; lin2022 ; newton2025a ; amesControlBarrierFunctions2019 ; clark2021 ; tan2006 ; schneeberger_sos_2023 ; loureiro2025 ). This section concludes with a comparison of the main characteristics of the two methods and some practical aspects such as initial guess generation for the sequential quadratic SOS approach.
5.1 Preliminaries
All benchmarks are implemented in CaΣoS Cunis2025 v1.0.0 and are available in the supplementary material DARUS-5677_2026 . In CaΣoS, the SOS problems are parameterized once at the beginning, which we refer as build process with its corresponding build time. The solve time is the total time spent in the iterative procedure. We make use of MOSEK mosek v11.1.8 as the underlying SDP solver. As we are not aware of existing benchmarks for nonconvex SOS problems, we make use of several SOS problems arising in systems and control theory with some modifications. Hence, detailed derivations of the SOS problem are omitted. The system dynamics and the considered polynomials are provided in the supplementary material DARUS-5677_2026 . If not differently stated, we assume CD is converged if $|f_{k}-f_{k-1}|\leq\epsilon_{\text{opt}}$, which is equivalent to the scaled termination criterion in (23). The maximum number of iterations is set to $N_{\text{max}}=100$. We compare the number of iterations, total computation time (total solve time + solver build time) and final cost for a given problem size. We denote the problem size by the number of system states $n$ and the number of constraints $n_{\text{con.}}$ (both linear and conic constraints) and decision variables $n_{\text{dec.}}$ of the underlying SDP. We report only the problem size of the sequential quadratic SOS approach, as it reflects the overall size. In contrast, CD involves multiple smaller subproblems, which may include redundant constraints or variables. A detailed breakdown is available in the supplementary material DARUS-5677_2026 .
All results were computed on a personal computer with Windows 11 and MATLAB 2023b running on an AMD Ryzen 9 5950X 16-Core Processor with 3.40 GHz and 128 GB RAM.
5.2 Region-of-Attraction Estimation
We consider the problem of computing an inner approximation for the region-of-attraction (ROA) of nonlinear closed-loop systems. We solve the standard ROA estimation problem
@@ 5.2 Region-of-Attraction Estimation
where $V$ is the sought Lyapunov function, $s$ is a SOS multiplier resulting from Theorem 2.1 and $f(\cdot)$ is the closed loop dynamic, $g_{0}$ is a pre-defined safe set which we make use of to find a less conservative solution by minimizing the squared $l_{2}$-norm $\|\cdot\|_{\mathbb{R}[x]}$ on the space of polynomials. A small cost value indicates large portions of the safe set are covered by the ROA function. Thus, such a cost function is a good choice for (state) constrained problems, to obtain less conservative results and is also used in subsequent benchmarks in a similar fashion. The stable level set is fixed to $\gamma=1$, but could be also a decision variable. Two examples are considered: The first example is taken from chakrabortyEtAl2011ii , where the nonlinear ROA analysis is used to assess the robustness properties of the F/A-18 flight control laws. The second example evaluates scalability and robustness in terms of the number of system states $n$, using an $N$-link (planar) robot arm.
5.2.1 F/A-18 Flight Control Laws
We take the seven state closed-loop dynamics from (chakrabortyEtAl2011ii, , Appendix B). For both methods, we compute an initial Lyapunov candidate by linearizing the nonlinear closed-loop dynamics. For the sequential quadratic SOS approach, we additionally choose the initial SOS multiplier $s=\sum_{i}x_{i}^{2}$. Due to infeasibility of CD at the first iteration, we solved the first subproblem via bisection. After the first iteration, we used the pre-defined level set. The results of the two approaches are summarized in Table 2.
The sequential quadratic SOS approach only needs 19 iterations until convergence, whereas coordinate descent did not converge within the pre-assigned maximum number of iterations. The total time (solver build + solve time) as stated in Table 2 is substantially less for the sequential quadratic SOS approach. The build time for the sequential quadratic SOS approach is $4.81\text{\,}\mathrm{s}$ opposed to only $1.15\text{\,}\mathrm{s}$ for coordinate descent, which includes several subproblems.
Table 2: Comparison of sequential quadratic SOS (SQ) and coordinate descent (CD) for the F/A-18 ROA estimation problem.
@@ 5.2.1 F/A-18 Flight Control Laws
5.2.2 N-link Robot Arm
Using an $N$-link (planar) robot arm, the problem scales to $n=2N$ states. For each $N$, the dynamics and a linear quadratic control law (LQR) are computed and a quadratic polynomial approximation for the closed-loop system is calculated. This preparation phase is provided in the supplementary material DARUS-5677_2026 . We solve the problem twice using different initial guesses. In the first case, both algorithms are initialized with the Lyapunov function of the linear system, with the sequential quadratic SOS algorithm additionally given a simple quadratic polynomial for the SOS multiplier. In the second, we make use of a clearly infeasible initial guess. A more detailed breakdown is given in Table 3 and  4.
Table 3: Comparison of sequential quadratic SOS (SQ) and coordinate descent (CD) for the N-link robotic arm ROA estimation problem using a good initial guess. indicates that CD was initially infeasible.
n Size Iterations [-] Time [s]
$_{\text{con.}} n$$_{\text{dec.}}$ SQ CD SQ CD
@@ 5.2.2 N-link Robot Arm
Good Initial Guess
[…]
Table 4: Comparison of sequential quadratic SOS (SQ) and coordinate descent (CD) for the N-link pendulum ROA estimation problem using a bad initial guess. indicates convergence to a feasible solution.
@@
$_{\text{con.}} n$$_{\text{dec.}}$ SQ CD SQ CD
44 $85$ $248$ $45$ $3.6$
66 $245$ $843$ $56$ $7.8$
88 $558$ $2136$ $73$ $30.9$
$10$ $1100$ $4535$ $11$ $16.7$
$12$ $1963$ $8544$ $12$ $52.3$
$14$ $3255$ 14 76314\,763 $11$ $150.2$
$16$ $5100$ 23 88823\,888 $12$ $454.2$
$18$ $7638$ 36 71136\,711 $16$ $1354.01$
2020 11 02511\,025 54 12054\,120 $15$ $3073.6$
$22$ 15 43315\,433 77 09977\,099 $14$ $9873.7$
$24$ 21 05021\,050 106 728106\,728 $15$ 22 245.922\,245.9
Bad Initial Guess
The second case considers the case with a bad initial guess. We initialize all decision variables with negative quadratic polynomials. Such polynomials are clearly infeasible. Coordinate descent is not able to start the computation because it needs a feasible initial guess. Thus, no results can be reported in this case. In contrast, the proposed sequential algorithm with its feasibility restoration is able to solve the problem to full optimality or, in a few cases, at least to a feasible solution [^1] That is, the thresholds for constraint violations are met, but the solution might be far from optimal. (indicated by in Table 4). This clearly shows the enhanced robustness of the proposed approach. The call of the feasibility restoration phase increases the number of iterations significantly for smaller number of states. For medium to large number of states, the number of iterations is only moderately increased. However, the average time per iteration is much larger compared to cases where the feasibility restoration is not called, which is expected.
[…]
Figure 2: Relative performance improvement of sequential quadratic SOS (SQ) with respect to coordinate-descent (CD) in terms of total computation time $T_{(\cdot)}$ and number of iterations $I_{(\cdot)}$.
Remark 2
It is unlikely that someone uses such an obviously infeasible initial guess and rather uses an initial guess similar to the first case. This is only made for demonstration purpose. It should be noted that an initial guess like in the first case does not guarantee convergence for coordinate descent as demonstrated in the previous example.
5.3 Control Synthesis for Nonlinear Systems
We consider the synthesis of a stabilizing control law for nonlinear systems subject to state and control constraints. Such problems are also relevant for computing terminal conditions that ensure asymptotic stability of nonlinear model predictive control. We consider the input constraint set as polyhedral set
@@ 5.3 Control Synthesis for Nonlinear Systems
with $H_{\mathcal{U}}\in\mathbb{R}^{p\times m}$, $p\in\mathbb{N}$, and $\mathbf{1}_{p}=(1,..,1)\in R^{p}$. We solve the nonconvex SOS problem
@@
where $V(\cdot)$ is a the sought Lyapunov function, $\beta>0$ is the level set (either fixed or an additional decision variable), $\hat{\kappa}$ is the sought polynomial control law, and $s(\cdot)$ are the SOS multipliers resulting from Theorem 2.1. Eqs. (25b) and (25c) are standard Lyapunov conditions whereas (25d) and (25e) enforce state and control constraints, respectively. We make use of a similar cost function as before.
Again two examples are considered: The first example considers the nonaffine nonlinear dynamics of an aircraft’s longitudinal motion with four states and two control inputs, which cannot be solved with coordinate descent in a direct manner. The second showcases the synthesis of a control law for control-affine satellite attitude dynamics.
5.3.1 Nonlinear Aircraft Dynamics
In this example, we make use of the four-state longitudinal motion of an aircraft provided in chakrabortyEtAl2011 . Compared to the previous examples, the control inputs enter nonaffinely into the dynamics and hence, the synthesis problem cannot be solved by coordinate descent in a direct manner. For synthesis, we must re-write the system dynamics into a control-affine form. This would be possible by augmenting the system dynamics with a zero dynamic of the controls, resulting in $n_{e}=n+m$ states and hence in a higher computational load. In contrast, with the sequential quadratic SOS approach we are able to directly solve this problem. In Table 5, we provide only results for the sequential quadratic SOS approach. In about $13\text{\,}\mathrm{s}$ one obtains a safe-by-design feedback law for the nonlinear system.
5.3.2 Spacecraft Attitude Control
The second example considers the control synthesis for a Hubble-like spaceborne telescope nurre1995 along with attitude, rate and control constraints. The details can be found in the supplementary material DARUS-5677_2026 . The six-state dynamics are comprised of angular rates and Modified-Rodrigues parameters, the latter of which describe the spacecraft attitude. The results are summarized in the second row of Table 5. Whereas the sequential quadratic SOS approach solves the problem in slightly above one minute to (local) optimality, the coordinate descent approach failed after six iterations (indicated by ), far from optimality.
Table 5: Comparison of sequential quadratic SOS (SQ) and coordinate descent (CD) for the constrained control synthesis problems. indicates failure.